1) Filename: MAE1090_HW7_1.m Write an m-file that determines the maximum value in a vector, x, and stores that maximum value as the variable y. Your m-file should work for both row and column vectors. For the purpose of testing and grading your m-file:

 DO NOT issue a clear command at the start of your m-file

 DO NOT define the vector x within your m-file. That vector should exist in the workspace prior to executing your m-file.

 Do not display y. This will be done by the script the grader uses to test your m-file.

 Variable names (x and y) must be exactly as specified (lower case).
Answer #1

MatLab Code

x + MAE1090_HW_7_1. m 1- y = x(1); WN for i= 2: length(x) if x(1) >y y = x(1); end end

Sample Output

>> X = (2,12.5, 54, 32,0.5, 67, 43, 51, 66]; >> MAE1090_HW_7_1 >> y >> x = [2,12.5,14,59,55,6,43,1,9,19]; >> MAE1090_HW_7_1